Daniel is paining the walls and ceiling of a cuboidal hall with length, breadth and height of 15 m, 10 m and 7 m respectively. From each can of paint 100 m^2of area is painted. How many cans of paint will she need to paint the room? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Length of wall, l = 15 m, Breadth of wall, b = 10 m and
Height of wall, h = 7 m
Total Surface area of classroom = lb+2(bh+hl )
= 15×10+2(10×7+7×15)
= 150+2(70+105)
= 150+350
= 500
Now, Required number of cans = Area of hall/Area of one can = 500/100 = 5
Therefore, 5 cans are required to paint the room.

Thanks